OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = - 3/4 x - 2?

OpenStudy (anonymous):

y = 4/2 x - 12 y = 3/4 x - 5 y = - 4/3 x + 20 y = - 3/4 x + 13 Which answer would it be then ?

OpenStudy (anonymous):

@AlwaystheBookworm

OpenStudy (anonymous):

y = - 3/4 x - 2 and the points (12, 4) Use y-y=m(x-x) and the inverse of the original slope. y-4=4/3(x-12) 3y-12=4x-48 3y=4x-48+12 3y=4x-36 y=4/3x-12

OpenStudy (dan815):